# Evaluating NeRFs for 3D Plant Geometry Reconstruction in Field Conditions

### v2

@@ 4. Discussion
Instant-NGP’s sampling strategy is built for speed, with an occupancy grid that helps skip irrelevant samples. This approach is great for real-time applications but can potentially miss subtle density variations, leading to a denser and less clear output if the grid isn’t capturing all the necessary detail. Nerfacto’s sampling strategy is more complex and layered, with multiple mechanisms in place to ensure that sampling is done more effectively in areas that greatly affect the visual output. The combination of pose refinement, piecewise sampling, proposal sampling, and an efficient density field leads to more accurate sampling, which in turn produces crisper images. In summary, the reason for Nerfacto’s superior crispness likely stems from its more refined and targeted approach to sampling, which concentrates computational efforts on the most visually impactful parts of the scene. In contrast, Instant-NGP’s faster but less targeted sampling may result in less clarity and more visual artifacts.
5. Conclusions
The findings of this research are important for precision agriculture, offering a non-destructive and efficient method for detailed plant analysis. The ability to accurately reconstruct plant structures in 3D without physically altering or harming the plant is a key advancement. This non-intrusive approach enables essential agricultural tasks such as growth monitoring, yield prediction, and early disease detection to be performed more effectively, without the need for physical sampling or destruction of the plants. Furthermore, this research extends its impact beyond immediate applications.
In addition to these methodological advancements, our study contributes valuable resources for ongoing research. We collected a comprehensive dataset comprising images and ground truth 3D LiDAR scans, which will be instrumental for further studies. This dataset not only facilitates the testing and validation of current models but also serves as a benchmark for the development and evaluation of future NeRF models in agricultural contexts. Moreover, we developed a robust framework for the evaluation of NeRF models, which can be adapted and utilized in future research. This framework is designed to assess various aspects of model performance, ensuring a comprehensive evaluation of new techniques as they emerge. Moreover, this groundwork sets the stage for our next focus.
Our research also provides detailed insights into the performance of each model, identifying their strengths and limitations in various scenarios. These insights and resources pave the way for future research at the nexus of NeRF technologies and agriculture. The detailed analysis and case studies presented in this paper will be invaluable for researchers and practitioners looking to leverage NeRF in agricultural applications, contributing to the development of more advanced, efficient, and sustainable agricultural practices.
Supplement
Table 6. Performance metrics of NeRFs reconstruction techniques - Scenario I
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 0.29 2.48 0.53 17.14 0.58 0.81 13
Instant-NGP 200 0.32 1.50 0.53 18.23 0.56 0.75 26
Instant-NGP 400 1.25 7.14 2.13 19.64 0.60 0.66 42
Instant-NGP 800 5.66 35.17 9.74 21.21 0.64 0.55 61
Instant-NGP 1000 3.93 27.61 6.88 20.71 0.57 0.59 76
Instant-NGP 5000 21.93 89.03 35.20 22.73 0.76 0.28 175
Instant-NGP 10000 25.98 92.59 40.57 23.20 0.79 0.22 297
Instant-NGP 20000 23.21 88.38 36.77 23.42 0.81 0.18 527
Instant-NGP 30000 24.66 90.62 38.77 23.41 0.81 0.17 756
NeRFacto 100 1.94 20.98 3.55 18.11 0.59 0.75 14
NeRFacto 200 7.72 42.84 13.08 19.50 0.57 0.65 27
NeRFacto 400 20.86 68.64 32.00 21.27 0.64 0.55 43
NeRFacto 800 39.48 80.26 52.92 22.33 0.67 0.45 64
NeRFacto 1000 41.35 82.54 55.09 22.20 0.66 0.44 79
NeRFacto 5000 66.43 92.51 77.33 22.30 0.73 0.19 430
NeRFacto 10000 70.04 93.94 80.25 22.34 0.74 0.15 564
NeRFacto 20000 73.32 94.51 82.58 22.35 0.74 0.13 1068
NeRFacto 30000 73.57 94.72 82.81 22.24 0.73 0.12 1938
TensoRF 100 0.43 2.27 0.72 13.44 0.54 0.82 13
TensoRF 200 0.65 4.47 1.13 13.55 0.52 0.82 25
TensoRF 400 1.18 8.76 2.07 13.51 0.51 0.81 40
TensoRF 800 1.86 14.22 3.29 13.10 0.50 0.79 61
TensoRF 1000 2.05 16.40 3.65 13.14 0.49 0.79 77
TensoRF 5000 6.63 36.57 11.22 13.59 0.52 0.70 420
TensoRF 10000 9.58 43.47 15.69 14.64 0.55 0.67 859
TensoRF 20000 9.51 43.19 15.59 14.68 0.55 0.67 1651
TensoRF 30000 9.58 43.34 15.69 14.69 0.55 0.66 1973
Table 7. Performance metrics of NeRFs reconstruction techniques - Scenario II
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 0.57 1.23 0.78 13.70 0.36 0.89 13
Instant-NGP 200 1.52 4.41 2.27 13.91 0.32 0.85 27
Instant-NGP 400 2.49 9.35 3.93 14.95 0.36 0.77 43
Instant-NGP 800 5.83 24.29 9.41 15.91 0.39 0.69 63
Instant-NGP 1000 5.72 22.80 9.14 15.68 0.32 0.69 78
Instant-NGP 5000 18.30 51.68 27.02 18.12 0.55 0.44 176
Instant-NGP 10000 20.86 55.25 30.28 18.67 0.59 0.38 296
Instant-NGP 20000 23.24 58.53 33.27 19.05 0.63 0.33 1023
Instant-NGP 30000 23.45 58.57 33.49 19.08 0.64 0.31 1886
NeRFacto 100 7.77 27.44 12.12 14.93 0.35 0.77 15
NeRFacto 200 16.42 37.77 22.89 15.51 0.32 0.70 28
NeRFacto 400 25.83 49.81 34.02 16.40 0.39 0.67 44
NeRFacto 800 35.97 56.86 44.07 17.10 0.43 0.60 64
NeRFacto 1000 36.22 57.80 44.53 16.96 0.40 0.61 79
NeRFacto 5000 58.64 70.87 64.18 18.69 0.59 0.37 404
NeRFacto 10000 61.31 74.12 67.11 18.91 0.62 0.31 749
NeRFacto 20000 63.68 76.21 69.38 19.00 0.64 0.27 988
NeRFacto 30000 64.47 76.80 70.10 18.93 0.64 0.25 1226
TensoRF 100 1.51 12.62 2.70 13.38 0.35 0.83 14
TensoRF 200 2.53 17.53 4.43 13.81 0.34 0.79 27
TensoRF 400 4.41 26.94 7.58 14.19 0.34 0.74 44
TensoRF 800 6.28 35.10 10.66 14.46 0.34 0.71 66
TensoRF 1000 7.09 35.61 11.82 14.51 0.33 0.70 82
TensoRF 5000 13.09 51.34 20.86 15.16 0.36 0.64 399
TensoRF 10000 20.51 55.03 29.88 15.51 0.42 0.57 840
TensoRF 20000 20.44 54.97 29.80 15.54 0.42 0.56 1709
TensoRF 30000 20.50 55.34 29.91 15.54 0.42 0.56 2607
Table 8. Performance metrics of NeRFs reconstruction techniques - Scenario III
Model Name Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time(s) $\downarrow$
Instant-NGP 100 1.61 17.61 2.95 15.31 0.29 0.91 16
Instant-NGP 200 1.29 8.19 2.23 14.66 0.27 0.84 30
Instant-NGP 400 2.45 19.22 4.34 15.65 0.29 0.82 46
Instant-NGP 800 2.81 18.90 4.89 15.91 0.29 0.79 67
Instant-NGP 1000 2.88 10.08 4.48 15.61 0.26 0.71 82
Instant-NGP 5000 6.47 40.28 11.15 17.39 0.37 0.61 186
Instant-NGP 10000 9.38 47.12 15.65 17.86 0.40 0.54 311
Instant-NGP 20000 11.62 54.26 19.13 18.27 0.43 0.48 548
Instant-NGP 30000 12.96 56.88 21.11 18.46 0.45 0.45 783
Instant-NGP 60000 15.06 59.55 24.04 18.54 0.47 0.40 1466
NeRFacto 100 4.26 35.99 7.61 15.65 0.29 0.84 13
NeRFacto 200 6.53 42.49 11.31 15.91 0.27 0.66 27
NeRFacto 400 10.48 50.45 17.35 16.44 0.29 0.69 43
NeRFacto 800 18.49 57.01 27.93 16.81 0.31 0.69 64
NeRFacto 1000 18.39 58.02 27.93 16.56 0.29 0.65 80
NeRFacto 5000 49.87 74.75 59.83 17.29 0.34 0.54 189
NeRFacto 10000 58.22 78.80 66.96 17.24 0.34 0.47 318
NeRFacto 20000 64.91 80.61 71.92 17.03 0.34 0.41 561
NeRFacto 30000 66.30 81.33 73.05 16.88 0.33 0.38 803
NeRFacto 60000 68.29 82.32 74.65 16.70 0.32 0.34 1499
TensoRF 100 1.48 12.65 2.65 15.05 0.29 0.96 17
TensoRF 200 3.05 27.40 5.48 15.26 0.29 0.90 31
TensoRF 400 5.69 43.45 10.06 15.63 0.29 0.83 47
TensoRF 800 8.92 49.59 15.13 15.96 0.29 0.78 68
TensoRF 1000 10.48 50.82 17.37 16.03 0.29 0.75 84
TensoRF 5000 28.74 70.83 40.89 16.84 0.34 0.64 208
TensoRF 10000 40.85 75.24 52.95 17.31 0.38 0.56 374
TensoRF 20000 40.82 75.38 52.96 17.33 0.39 0.55 697
TensoRF 30000 40.80 75.26 52.92 17.34 0.39 0.55 1018
TensoRF 60000 40.95 75.62 53.13 17.32 0.39 0.55 1965
(a) 200 Iterations
(b) 400 Iterations
(c) 800 Iterations
(d) 5000 Iterations
(e) 10000 Iterations
(f) 30000 Iterations
Figure 22. Precision Over Training Iterations for NeRFacto - Scenario II
Table 9. Performance metrics of MipNeRF (failed) reconstruction - Scenario II
Iterations Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time (s) $\downarrow$
100 0.47 7.57 0.89 11.01 0.34 0.91 26
200 0.63 10.67 1.19 11.24 0.34 0.93 52
400 0.69 12.1 1.3 11.07 0.34 0.92 86
800 0.48 7.68 0.91 9.88 0.33 0.94 158
1000 0.44 7.19 0.82 9.65 0.32 0.94 192
5000 0.35 5.41 0.66 9.31 0.31 0.94 687
10000 0.42 6.29 0.79 8.88 0.29 0.9 1303
20000 0.34 5.3 0.65 8.78 0.28 0.89 2508
30000 0.35 5.71 0.67 9 0.27 0.88 3856

### v3

@@ 4   Discussion
Instant-NGP’s sampling strategy is built for speed, with an occupancy grid that helps skip irrelevant samples. This approach is great for real-time applications but can potentially miss subtle density variations, leading to a denser and less clear output if the grid isn’t capturing all the necessary detail. Nerfacto’s sampling strategy is more complex and layered, with multiple mechanisms in place to ensure that sampling is done more effectively in areas that greatly affect the visual output. The combination of pose refinement, piecewise sampling, proposal sampling, and an efficient density field leads to more accurate sampling, which in turn produces crisper images. In summary, the reason for Nerfacto’s better reconstruction likely stems from its more refined and targeted approach to sampling, which concentrates computational efforts on the most visually impactful parts of the scene. In contrast, Instant-NGP’s faster but less targeted sampling may result in less clarity and more visual artifacts.
Finally, to retrieve the scale of the 3D reconstruction in the absence of reference point cloud data, a known scale can be placed on the ground during data collection. The exported point cloud can then be proportionally scaled based on this reference scale, which allows the size of the reconstructed plant to be calibrated to match its real-world dimensions. In order to show the practicality of this approach, we placed a 3D printed sphere of known diameter for the plant in Scenario I and captured the images. We then go through our pipeline of NeRF reconstruction, and instead of registering and scaling the scene to the ground truth LiDAR data, we scale it to the known sphere size. We then measured the height of the plant in this scenario. We find that by using this approach, the error in the height of the plant is within 1%. We provide additional details of this experiment in the Supplement. We note that this is a preliminary result, and more detailed studies need to be performed in the future on extracting the correct scale from NeRF reconstructions.
5   Conclusions
The findings of this research underscore the value of NeRFs as a non-destructive approach for 3D plant reconstruction in precision agriculture. Our methodology more effectively facilitates critical agricultural tasks, such as growth monitoring, yield prediction, and early disease detection from accurate reconstruction of plant structures. Our comparative analysis, which benchmarks different NeRF models against ground truth data, highlights the method’s efficiency, achieving a 74.65% F1 score within 30 minutes of GPU training. Introducing an early stopping algorithm based on LPIPS further enhances this process, reducing training time by 61.1% while limiting the average F1 score loss to just 7.4%.
Additionally, our work provides a comprehensive dataset and an evaluation framework, aiding the validation of current models and serving as a foundation for developing future NeRF applications in agriculture. The detailed insights into model performance across varied scenarios, coupled with the early stopping case study, offer practical guidance for 3D reconstruction using NeRFs. This research supports the advancement of non-intrusive agricultural technologies and also sets a baseline for future work at the intersection of NeRF technologies and agriculture, aiming to improve efficiency and accuracy in plant phenotyping and breeding.
Acknowledgements
This work was supported in part by the Plant Science Institute at Iowa State University, the National Science Foundation under grant number OAC:1750865, and the National Institute of Food and Agriculture (USDA-NIFA) as part of the AI Institute for Resilient Agriculture (AIIRA), grant number 2021-67021-35329.
Supplementary Materials
Figures S1 to S9.
Tables S1 to S5.
Author Contributions
Muhammad Arbab Arshad: Methodology, Formal analysis, Software, Writing–Original Draft, Data Curation, Visualization.
Talukder Jubery: Data Curation, Supervision, Writing–Review & Editing.
James Afful: Data Curation.
Anushrut Jignasu: Data Curation.
Aditya Balu: Methodology, Supervision, Writing–Review & Editing.
Baskar Ganapathysubramanian: Conceptualization, Methodology, Supervision, Project administration.
Soumik Sarkar: Conceptualization, Methodology, Project administration, Funding acquisition.
Adarsh Krishnamurthy: Conceptualization, Methodology, Supervision, Writing–Review & Editing, Project administration, Funding acquisition.
Data Availability
The data for the four Scenarios including raw images and the point cloud will be made available online. The GIT repository containing the different implementations of the NeRF code will also be made public.